At graduation, you have a student loan of $19,000. The loan has a fixed annual interest rate of 3.4%, compounded monthly, for 10 years. What will your monthly payments be? (Enter numeric answers to 2 decimal places.)
Found 2 solutions by ikleyn, math_tutor2020:
Answer by ikleyn(52781)

.
At graduation, you have a student loan of $19,000.
The loan has a fixed annual interest rate of 3.4%, compounded monthly, for 10 years.
What will your monthly payments be? (Enter numeric answers to 2 decimal places.)
~~~~~~~~~~~~~~~~~~~~~


Go to web-site

https://www.calculatorsoup.com/calculators/financial/loan-calculator.php

Find there free of charge online calculator, specially developed
to make this kind of calculations for """students""" homework
with complete explanations.

Print your input data in calculator and get your answer immediately.


In parallel, learn the relevant formulas and calculation technique from the accompanying text.



Answer by math_tutor2020(3817)


Answer: $186.99

Work Shown:

The formula to use is
P = (L*i)/( 1-(1+i)^(-n) )
where,
P = monthly payment
L = loan amount
i = monthly interest rate in decimal form
n = number of months

In this case
L = 19000
i = 0.034/12 = 0.0028333 approximately
n = 10*12 = 120 months

Then,
P = (L*i)/( 1-(1+i)^(-n) )
P = (19000*0.0028333)/( 1-(1+0.0028333)^(-120) )
P = 186.994047599858
P = 186.99

The answer can be confirmed with a calculator such as this one
https://www.calculator.net/loan-calculator.html